OpenStudy (anonymous):

Thirty is 12% of what number?

OpenStudy (anonymous):

Set up a proportion. 30/x = 12/100 Solve for x by cross-multiplying.

OpenStudy (the_fizicx99):

12/100 = .12 30/.12 = 250 250*.12 = 30 <-- making sure

OpenStudy (anonymous):

250 after solving for x by cross - multiplying :) ..